OpenStudy (anonymous):

15. Which represents the solution to the absolute value equation? 2|x+2|-1=5 (1 point) x=4 or x=-8 x=1 or x=-1 x=1 or x=-5 x=5 or x=-5

OpenStudy (texaschic101):

2 | x + 2 | - 1 = 5 --- add 1 to both sides 2 | x + 2 | = 5 + 1 -- combine like terms 2 | x + 2 | = 6 -- divide both sides by 2 | x + 2 | = 3 x + 2 = 3 x = 3 - 2 x = 1 -(x + 2) = 3 -x - 2 = 3 -x = 3 + 2 -x = 5 x = -5 solutions are : x = 1 and x = -5

OpenStudy (anonymous):

Thank you for the fast reply. Makes sense now!

OpenStudy (texaschic101):

no problem...glad to help